OpenStudy (anonymous):

Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (2,6) and (-4, 8). (4 points) Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form. (3 points)

OpenStudy (campbell_st):

step 1 find the slope of the line using the ordered pairs (2, 6) and (-4, 8) can you do that..?

OpenStudy (anonymous):

whats the formula again?

OpenStudy (campbell_st):

its simply m = rise/run or \[m=\frac{y_{2} - y_{1}}{x_{2} - x_{1}}\]

OpenStudy (anonymous):

so -3 right?

OpenStudy (campbell_st):

don't that that quite works...

OpenStudy (anonymous):

whoops. flipped it. -2/3

OpenStudy (campbell_st):

thats still not quite right (8 - 6)/(-4 -2) = -2/6 or - 1/3

OpenStudy (campbell_st):

now given a point (2, 6) and a slope m = -1/3 you need to find the equation of the line...

OpenStudy (campbell_st):

any ideas on that to do...there are lots of choices...

OpenStudy (anonymous):

plug it into the formula Y=MX+B right?

OpenStudy (campbell_st):

ok... what you you get

OpenStudy (anonymous):

what would B be?

OpenStudy (campbell_st):

ok so you have the equation y = -1/3 x + b use the point (2, 6), substitute x = 2 and y = 6 into the equation to find b

OpenStudy (anonymous):

B=9

OpenStudy (campbell_st):

wow... 6 = -1/3 * (2) + b 6 = -2/3 + b b = 20/3 does that make sense...?

OpenStudy (anonymous):

I see what i did wrong yeah

OpenStudy (campbell_st):

so the equation of the line in slope intercept form is y = -1/3 x + 22/3 I'll let you do the standard form

OpenStudy (anonymous):

-1x -3y + 20